In a neon discharge tube, 2.9 x 1018 Ne+ ions move to the right per second, while 1.2 x 1018 electrons move to the left per second. The current in the discharge tube is?

Grade:12th pass

2 Answers

Arun
5 years ago
Dear Fatima
 
Net current i = (2.9 * 10^18 + 1.2 * 10^18)* e = 4.1 * 10^18 * 1.6 * 10^-19
 
 =  0.656 amp towards right
Rishi Sharma
3 years ago
Dear Student,
Please find below the solution to your problem.

Net current i = (2.9 * 10^18 + 1.2 * 10^18)* e = 4.1 * 10^18 * 1.6 * 10^-19
= 0.656 amp towards right

Thanks and Regards